4. An FCIO3 molecule can adoptfour orientations in the solid with negligible different in energy. What is the residual molar entropy?
Answer #1

The entropy which the crystal of a substance has at T=0 is called the residual entropy.

Entropy is denoted as S(0)

S(0) = k * In 4^N = k*N*In 4   (as 4 orientations)

For the sake of simplicity we will assume that only 1 mol of the sample is taken, therefore, Kn= R .

Then

S(0) = R In 4 = 2.303 R log 4 = 11.5 J/K.mol